OpenStudy (anonymous):

A student writes an incorrect step while checking if the sum of the measures of the two remote interior angles of triangle ABC below is equal to the measure of the exterior angle. Step 1: m∠m + m∠n + m∠o = 180 degrees (sum of angles of a triangle) Step 2: m∠p + m∠o = 180 degrees (adjacent supplementary angles) Step 3: Therefore, m∠m + m∠n + m∠o = m∠n + m∠p Step 4: So, m∠m + m∠n = m∠p

OpenStudy (anonymous):

In which step did the student first make a mistake and how can it be corrected? Step 1; it should be m∠m + m∠n + m∠o = 90 degrees (corresponding angles) Step 1; it should be m∠m + m∠n + m∠o = 90 degrees (adjacent angles) Step 3; it should be m∠m + m∠n + m∠o = m∠o + m∠p Step 2; it should be m∠o − m∠p = 90 degrees (complementary angles)

OpenStudy (anonymous):

@misty1212 @amistre64

OpenStudy (anonymous):

@texaschic101 @Nnesha

OpenStudy (anonymous):

@TheSmartOne @sleepyjess

OpenStudy (anonymous):

@StudyGurl14 @robtobey

OpenStudy (studygurl14):

It would be step 3, because <n and <p are not supplementary

OpenStudy (anonymous):

ohh ok thanks cuz ive been on this question like forever ;)